# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting

### v1

@@ Abstract.
Reconstructing textured meshes from colored point clouds is an important but challenging task in 3D graphics and vision. Most existing methods predict colors as implicit functions in 3D or UV space, suffering from blurry textures or the lack of generalization capability. Addressing this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud. It produces meshes with enhanced fidelity and clarity by 2D image inpainting, taking advantage of the mature techniques and massive data of 2D vision. Specifically, we first project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model. Next, we design a novel Non-Border-First strategy to unproject the colors of the inpainted dense images back to 3D space, thus obtaining the final textured mesh. In this way, our PointDreamer works in a zero-shot manner, requiring no extra training. Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068). Code at: https://github.com/YuQiao0303/PointDreamer.
@@
3D Textured Mesh, Colored Point Cloud Reconstruction, Diffusion Models, Image Inpainting
submissionid: ccs: Computing methodologies Reconstructionccs: Computing methodologies 3D imagingccs: Computing methodologies Shape representationsccs: Computing methodologies Appearance and texture representationsccs: Computing methodologies Shape inference
Figure 1. PointDreamer is a zero-shot framework to reconstruct 3D textured meshes from colored point clouds. The key idea is to utilize 2D diffusion priors to inpaint the projected 2D multi-view sparse images to dense ones, which are finally unprojected back to 3D to produce mesh textures. PointDreamer produces textures with enhanced clarity compared to various competitors, such as Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013), Neural Kernel Surface Reconstruction (NKSR) (Huang et al., 2023), and Texture Field (Oechsle et al., 2019).
[…]
teaser
1. Introduction
@@
3D textured mesh reconstruction from colored point cloud is a pivotal task in computer vision, graphics and robotics. It enables the transition from sparse and unstructured point clouds to connected and visually-appealing meshes, facilitating applications like digital twin (Tao et al., 2022), metaverse (Wang et al., 2022a), cultural heritage preservation (Wang et al., 2024), etc.
With the great success achieved in geometry reconstruction by learning a 3D implicit field like signed distance field (SDF)  (Park et al., 2019), occupancy field (Mescheder et al., 2019; Chen and Zhang, 2019), etc., researchers propose to learn a 3D texture field (Oechsle et al., 2019) representing color information in 3D space. Nevertheless, such approaches often yield blurry appearances (Tang et al., 2023a). In addition, they necessitate 3D data for training, which is notoriously challenging to acquire, and the limited training data further hinder the models’ generalization capability to unseen objects.
While the 3D textured mesh reconstruction task remains challenging, 2D vision has recently witnessed great prosperity in the field of image super-resolution and image generation, particularly thanks to the emergence of diffusion models (Ho et al., 2020). Compared to 3D vision, 2D vision is more mature, and the abundance of 2D data offers a robust prior, which has been proven crucial for enhancing the performance and generalization ability of deep learning models (Sarker, 2021; Naveed et al., 2023). Inspired by this, numerous studies (Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023b; Lin et al., 2023; Qian et al., 2024) seek to adopt 2D diffusion models as a prior to guide 3D generation , mostly by optimizing a score distillation sampling (SDS) (Poole et al., 2022) loss or reconstruction loss between the generated 3D models’ rendered images and 2D diffusion models’ output images. However, they focus on text/image/un-conditioned 3D generation, leaving point-cloud-based reconstruction unexplored. Besides, the colors or textures are still learned or optimized in 3D space (as a texture field or radiance field), yielding blurry appearances (Tang et al., 2023a).
Going beyond existing approaches, our intuition is that, reconstructing 3D textured meshes from colored point clouds is analogous to inpainting 2D sparse images by filling empty pixels. Both tasks focus on somehow dreaming the missing areas of the sparse input, to achieve a more complete and coherent representation. Intuitively, dreaming in 2D space is much easier than in 3D space, especially with powerful tools like diffusion models. This inspires us to convert the dreaming process into 2D space to enhance reconstruction quality especially clarity, and then reconvert the results back to 3D. Specifically, unlike existing works (Gao et al., 2022; Sun et al., 2023; Liu et al., 2024; Poole et al., 2022) that learns per-point colors as implicit functions in 3D space, we predict per-pixel colors in 2D image space by 2D image inpainting, and then directly unproject the high-quality results back to 3D space, so as to generate clear textures for meshes.
Guided by this idea, we design PointDreamer, a novel zero-shot framework for reconstructing high-quality textured meshes from colored point clouds, as shown in Figure 1. With only a set of pre-defined camera parameters from different viewpoints, an off-the-shelf 2D diffusion model, and a geometry extraction module, our method starts from projecting the input point cloud into 2D to generate multi-view sparse images. The 2D diffusion model is then adopted to inpaint the sparse images into dense ones. Finally, we unproject the colors of dense images back to obtain the 3D textured mesh by a carefully-designed Non-Border-First unprojection strategy. It prioritizes non-border areas during unprojection, to circumvent inaccurate color prediction at occlusion borders and thus avoid artifacts. Overall, we list our contributions as follows:
@@
  • We propose PointDreamer, a zero-shot framework for 3D textured mesh reconstruction from colored point cloud. By novelly transforming this 3D reconstruction task into 2D inpainting, PointDreamer achieves high-fidelity appearance and strong generalization ability while requiring no extra training, and pioneers in utilizing 2D diffusion models for 3D point cloud reconstruction.
@@
  • We design Non-Border First (NBF) Unprojection, a novel unprojection strategy for 2D-to-3D re-conversion that effectively avoids artifacts caused by inaccurate color prediction near occlusion borders.
@@
  • Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods both quantitatively and qualitatively.
2. Related Work
2.1. Textured Mesh Reconstruction from Colored Point Cloud
While many works (Mescheder et al., 2019; Peng et al., 2020; Tang et al., 2021; Boulch and Marlet, 2022) focus on reconstructing surfaces from non-colored point clouds, textured mesh reconstruction from colored point cloud remains a field to be further explored. Traditionally, (Screened) Poisson Surface Reconstruction (SPR) (Kazhdan et al., 2006; Kazhdan and Hoppe, 2013) reconstructs textured meshes by solving the Poisson equation, but it relies on accurate per-point normals as input. In contrast, DHSP3D (Wei et al., 2021) optimizes a MeshCNN in 3D space and XYZ/RGB maps in 2D UV space by self-supervision, but it does not support arbitrary topology. Recently, a more common practice (Gupta et al., 2023; Huang et al., 2023) is to learn two 3D implicit fields or hyper-representations for geometry and texture, respectively. For geometry reconstruction, Signed Distance Field (SDF) (Gupta et al., 2023), Occupancy Field (Chen et al., 2022), Neural Kernel Field (Huang et al., 2023), Deep Marching Tetrahedra (DMTet) (Shen et al., 2021; Gupta et al., 2023), etc. are optional shape representations to be learned or optimized. For texture reconstruction, a texture field (Oechsle et al., 2019) consisting of separated functions for red, green and blue channels is generally fitted, supervised by MAE loss for per-point color (Chen et al., 2022; Huang et al., 2023) or reconstruction loss (Gupta et al., 2023) for images from differentiable rendering. Unfortunately, such methods without powerful prior often suffer from relatively blurry texture (Kerbl et al., 2023). To address this, we are the first, to the best of our knowledge, to perform the 3D reconstruction task by 2D inpainting to take advantage of existing high-performance 2D vision techniques.
2.2. Representations of Colored 3D Objects
Researchers have explored various ways to represent 3D objects or scenes with colors. To begin with, CMR (Kanazawa et al., 2018) and its followers (Li et al., 2020; Goel et al., 2020; Hu et al., 2021; Zhang et al., 2022) regard a shape as the deformation of a sphere with fixed UV mapping, thus only supporting spherical topology. Unlike them, NeRF variants (Mildenhall et al., 2020; Barron et al., 2021; Müller et al., 2022) or DMTet (Shen et al., 2021) with texture field (Oechsle et al., 2019) have shown their advantages in representing power and generation quality  (Gao et al., 2022; Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023c; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) thanks to the techniques of volumetric rendering and differentiable rendering (Laine et al., 2020). Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has drawn considerable attention where 3D scenes are represented as a set of 3D Gaussians. However, as mentioned before, learning or optimizing colors in 3D space often results in blurry textures (Tang et al., 2023a). In contrast, we directly predict high-fidelity colors in 2D image space by utilizing the power of a 2D pre-trained diffusion model, which can be further unprojected to 3D meshes to obtain clearer textures.
2.3. Diffusion Models for 3D Shape Generation
@@
With the success of 2D diffusion models, researchers have explored extending applications of diffusion models to 3D field. An intuitive idea is to apply the diffusion process on 3D representations, such as a DMTet grid (Liu et al., 2023a), occupancy or SDF field (Zheng et al., 2023), or 3D feature tri-plane (Gupta et al., 2023). To fully capitalize on the abundant availability of 2D data, more researchers seek to directly “lift” results of 2D diffusion models to 3D. Most of these works (Gao et al., 2022; Poole et al., 2022; Lin et al., 2023; Qian et al., 2024; Melas-Kyriazi et al., 2023; Liu et al., 2023c; Shi et al., 2023; Liu et al., 2023d, b; Tang et al., 2023c, b; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) optimize a NeRF, a DMTet or 3D Gaussianss by SDS (Poole et al., 2022) or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model (e.g. Stable-Diffusion). View-conditioned diffusion models (Liu et al., 2023c) or multi-view diffusion models (Szymanowicz et al., 2023; Liu et al., 2024) are developed by fine-tuning existing models for generating multi-view-consistent images. Most of these works still learn or optimize colors in 3D space. They tend to be time-consuming during optimization, and suffer from relatively blurry results in unseen views. Instead of optimizing a 3D representation from scratch to follow the images generated by 2D diffusion models, we propose to directly adopt colors of the 2D images, by unprojecting them back to 3D, for higher-quality texture reconstruction.
3. Method
@@
Given a 3D point cloud with per-point XYZ coordinates and RGB colors, our goal is to reconstruct its associated textured mesh. Unlike existing works that learn or optimize colors in 3D space (Oechsle et al., 2019; Poole et al., 2022) or UV space (Wei et al., 2021; Chen et al., 2022), we propose PointDreamer, a novel approach by conducting color prediction in 2D image space to leverage the powerful 2D diffusion priors. Figure 2 presents our designed pipeline with four steps. First, we employ a geometry extraction module to reconstruct an untextured mesh from the input point cloud. Next, with a set of fixed viewpoints, we perform 3D-to-2D conversion by projecting the input point cloud into 2D, producing sparse multi-view images. After that, we conduct color prediction in 2D space by inpainting the empty pixels in the sparse images to form dense ones based on a pre-trained 2D diffusion model (Dhariwal and Nichol, 2021). Finally, we propose a novel Non-Border-First strategy to convert the 2D results back to 3D by unprojecting the colors in the dense images to the untextured mesh to produce the desired textured mesh. Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training on 3D datasets.
Figure 2. Pipeline of our method. Given a colored point cloud, we first reconstruct its untextured mesh. Then, the colored point cloud is projected to 2D to produce sparse multi-view images, which are further inpainted to form the dense ones. Finally, we unproject the results back to 3D to output the associated textured mesh.
@@
Figure 3. Illustration of hidden point removal, without which the results would be messed up.
@@
Figure 4. Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh. Middle Left: An naive unprojection implementation. For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$. Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with inaccurate color since it’s near an occlusion border, producing an artifact. Bottom Right: our proposed Non-Border-First Unprojection. Observing the correspondence between occlusion border in 2D image space and visibility in UV space (both marked by sky blue dashed lines), we prioritize non-border-areas in UV space to avoid such artifacts.
3.1. Geometry Extraction: Point to Surface
@@
The first step in PointDreamer is to reconstruct an untextured mesh from the input point cloud. Since color information does not need to be considered in this step, many existing point-to-surface methods (Mescheder et al., 2019; Peng et al., 2020; Shen et al., 2021; Boulch and Marlet, 2022) can be employed. In our implementation, we directly adopt the state-of-the-art POCO (Boulch and Marlet, 2022) considering its high performance. Section 4 provides ablation experiments by using different geometry reconstruction methods.
3.2. Projection and Inpainting
As shown in Figure 2, we design a projection module for 3D-to-2D conversion, followed by an inpainting module for 2D color prediction by inpainting sparse multi-view images into dense ones.
3.2.1. Projection: 3D to 2D

### v2

@@ Abstract
Reconstructing textured meshes from colored point clouds is an important but challenging task. Most existing methods yield blurry-looking textures or rely on 3D training data that are hard to acquire. Regarding this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud via diffusion-based 2D inpainting. Specifically, we first reconstruct an untextured mesh. Next, we project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model. After that, we unproject the colors of the inpainted dense images onto the untextured mesh, thus obtaining the final textured mesh. This project-inpaint-unproject pipeline bridges the gap between 3D point clouds and 2D diffusion models for the first time. Thanks to the powerful 2D diffusion model pre-trained on extensive 2D data, PointDreamer reconstructs clear, high-quality textures with high robustness to sparse or noisy input. Also, it’s zero-shot requiring no extra training. In addition, we design Non-Border-First unprojection strategy to address the border-area inconsistency issue, which is less explored but commonly-occurred in methods that generate 3D textures from multiview images. Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068). Code at: https://github.com/YuQiao0303/PointDreamer.
Index Terms:
@@
Point Cloud Reconstruction, Texture Mapping, 2D Diffusion, Image Inpainting.
1 Introduction
@@
3D point clouds and meshes are two of the most fundamental 3D modalities in computer vision, graphics, and robotics. Reconstructing 3D textured meshes from colored point clouds (abbr. colored-PC-to-mesh) is a pivotal task, with extensive applications like digital twin [1], metaverse [2], cultural heritage preservation [3], etc. It requires the accurate reconstruction of both geometry and colors, to transform sparse and unstructured data into coherent and visually appealing models.
While recent methods [4, 5, 6] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds. Unlike images that provide dense pixel-level representations, the inherent sparse nature of point clouds makes it difficult to reconstruct dense colors. Regarding this, existing methods either blend input points’ colors by interpolation [7] or overfitting [8, 9], or train color prediction networks [10, 11] with 3D datasets. However, they often yield blurring-looking textures; see SPR [7], NKSR [9], and Texture Field [12, 10] in Figure 1. Also, 3D training data are notoriously challenging to acquire.
Figure 1: PointDreamer is a zero-shot framework to reconstruct 3D textured meshes from colored point clouds. It produces textures with higher quality compared to baseline methods. Its specially designed NBF unprojection strategy effectively addresses the border-area artifact issue.
While the colored-PC-to-mesh task remains challenging, 2D image generation has recently thrived with impressive plausibility, level of detail, and generalization capability. Therefore, many works [13, 14] leverage 2D diffusion models for 3D generation. Since the scale of existing 2D datasets far exceeds that of 3D datasets, leveraging diffusion models pre-trained on extensive 2D datasets can also significantly alleviate the reliance on 3D datasets. This inspires us to adopt 2D diffusion models for the colored-PC-to-mesh task, to address the low-quality color reconstruction issue.
However, it is not trivial to implement this idea: (1) Existing 2D-diffusion-for-3D methods typically focus on 3D generation conditioned on text [15, 16] or images [17, 18], leaving point-cloud-based reconstruction unexplored. (2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [19] by predicting colors in 3D space as a texture [20, 16]/radiance [21, 22, 23] field or 3D Gaussians [19, 18]. How to leverage 2D reference images to generate high-quality textures requires more attention. (3) Intermediate 2D multiview images generated by diffusion models often have inconsistencies with each other or with the reconstructed 3D geometry at occlusion border areas. When one part of an object occludes another, the generated 2D images and 3D geometry should predict highly consistent occlusion borders. Failing to do so would cause artifacts; see “Ours w/o/ NBF” in Figure 1, with significant artifacts near the border of the pillow and the chair.
We address these challenges one by one. First, to bridge the gap between 2D diffusion models and 3D point clouds, our intuition is that, reconstructing 3D textured meshes from colored point clouds is analogous to inpainting 2D sparse images by filling empty pixels. They both aim to somehow dreaming the missing areas of the sparse input, to achieve a more complete and coherent representation. Luckily, 2D diffusion-based inpainting methods [24] excel even with mask ratios over 80%, making them robust to sparse input points. Trained on large-scale 2D datasets, these models generalize well across domains and require no additional training or fine-tuning, enabling zero-shot point cloud reconstruction. Second, to address the blurring issue, we propose to directly unproject the generated multiview images onto the 3D mesh, instead of predicting colors in 3D [10, 9] or UV [8] spaces like existing methods. Third, to address the occlusion-border inconsistency issue, we design a “Non-Border-First” (NBF) unprojection strategy. It detected border areas by leveraging the correspondence of occlusion border in 2D image space and invisibility border in UV space, and then prioritizes non-border views’ images during unprojection, thus avoiding artifacts.
Putting everything together, we propose PointDreamer, a novel zero-shot framework to reconstruct high-quality textured meshes from colored point clouds, as shown in Figure 1. First, we extract geometry (untextured mesh) from the point cloud by an existing method POCO [4]. Second, we project the point cloud into 2D space from a fixed set of viewpoints, producing multiview sparse images. Third, we inpaint the empty pixels with an off-the-shelf 2D diffusion model, forming dense images. Finally, we directly unproject the 2D images onto the mesh. Unlike most existing methods that predict colors in 3D [10, 11] or UV spaces [8], directly adopting colors of 2D diffusion models’ output high-quality 2D images leads to clear textures. In particular, our designed NBF unprojection strategy effectively avoids artifacts caused by inconsistencies in occlusion border regions.
Overall, we list our contributions below:
@@
  • We propose PointDreamer, a zero-shot SoTA framework that fills the gap in the less-explored but important task of 3D textured mesh reconstruction from colored point cloud, pioneering in utilizing 2D diffusion models for 3D point cloud reconstruction.
  • PointDreamer reconstructs high-quality and clear-looking textures for meshes, avoiding the common blurring issue. PointDreamer is also zero-shot. While requiring no extra training, it shows high generalization capability: robust to noisy, sparse, or even incomplete input points, benefiting from the power of 2D diffusion models.
@@
  • We design Non-Border-First (NBF) Unprojection, a novel unprojection strategy to apply colors of 2D multiview images onto 3D meshes. It effectively avoids artifacts caused by inconsistent prediction near occlusion borders in any multiview-images-to-3D methods.
@@
Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods quantitatively and qualitatively.
2 Related Work
2.1 Textured Mesh Reconstruction from Colored Point Cloud
Existing methods can be categorized based on their reliance on training data. Non-data-driven methods rely solely on the input point cloud. For instance, (Screened) Poisson Surface Reconstruction (PSR, SPSR) [25, 7] reconstructs geometry by solving Poisson equations, and many 3D processing tools [26, 27] extend it with color support via linearly blending input point colors. DHSP3D [8] iteratively optimizes a MeshCNN in 3D space and XYZ/RGB maps in UV space by self-supervision. On the other hand, data-driven methods train neural networks on 3D datasets to generate textured meshes. For example, 3DGen [10] trains a triplane variational autoencoder to predict signed distance field and color values for each 3D tetrahedra grid vertex, and extracts textured meshes by marching tetrahedra [28]. Hybrid methods employ different strategies for geometry and texture reconstruction. NKSR [9] adopts data-driven geometry reconstruction by learning neural kernel fields, while using non-data-driven color reconstruction (in its supplementary file), by optimizing a 3D textured field [12] to overfit the input points’ colors. ColorMesh [11] employs non-data-driven geometry reconstruction via graph cuts, paired with a data-driven texture network to inpaint colors in 2D image space. However, its texture network is an encoder-decoder CNN trained with limited data. This may limit its performance and generalization capability, compared to our adopted zero-shot 2D-diffusion-based inpainting approach trained with extensive 2D data.
2.2 2D Diffusion Models for 3D Generation
@@ 2.1 Textured Mesh Reconstruction from Colored Point Cloud
To capitalize on the abundant availability of 2D data and the high performance of 2D diffusion models, researchers seek to “lift” results of 2D diffusion models to 3D. Most of these works optimize or learn a NeRF [22, 23], a DMTet with texture field [29, 16], or 3D Gaussianss by Score Distillation Sampling [13] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model. View-conditioned diffusion models [22] or multi-view diffusion models [30, 14, 31] are developed by fine-tuning existing diffusion models for generating multi-view-consistent images. Most of these works focus on 3D generation from text or images instead of point clouds. They tend to generate blurry textures, and may suffer from artifacts in occlusion border areas.
2.3 Border-Area-Inconsistency Issue in 3D Texturing
Mesh texture reconstruction based on real-scanned or AI-generated multiview images is widely used in applications such as 3D reconstruction and generation [32]. Due to scanning errors and imperfections in generative models, inconsistencies often arise between different views’ images or between images and the mesh, especially around occlusion borders where one part of the object occludes another. However, this issue remains largely unexplored. Most texture mapping methods [33, 34, 35, 36] design complex optimization objectives and strategies for camera extrinsics, etc. They typically focus on issues like blurring, ghosting, and seams, leaving border-area-inconsistency unexplored. Also, they predominantly rely on RGBD data, and thus are unsuitable for scenarios without dense depth data like point cloud reconstruction. On the other hand, most 3D generation approaches directly fuse multiview images’ colors for mesh texturing. Specifically, they use these images as supervision during optimization [32], or directly compute image colors’ weighted sum as the mesh’s colors [37]. To address the border-area-inconsistency issue, we design a Non-Border-First unprojection strategy, and it can be used in any method that unprojects multiview images to a mesh.
3 Method
@@
Given a 3D point cloud with per-point XYZ coordinates and RGB colors, our goal is to reconstruct its associated textured mesh. To avoid the blurring effect of existing works, we predict colors in 2D space by diffusion-based 2D inpainting, leveraging the powerful 2D diffusion priors. Figure 1 presents our designed pipeline with four steps. (1) We employ a geometry extraction module to reconstruct an untextured mesh from the input point cloud. (2) With a set of fixed viewpoints, we perform 3D-to-2D conversion by projecting the input point cloud into 2D, producing sparse multi-view images. (3) We conduct color prediction in 2D space by inpainting the empty pixels in the sparse images to form dense ones based on a pre-trained 2D diffusion model [38]. (4) We propose a novel Non-Border-First strategy to convert the 2D results back to 3D by unprojecting the colors in the dense images to the untextured mesh to produce the desired textured mesh.
Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training. It takes only about 61s and 30s for PointDreamer’s DDNM and nearest/linear interpolation variants (see Figure 7 (b)) to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [22] (~20 m), DreamGaussin [19] (~2 m), Wonder3D [32] (2~3 m), LGM [18] (~65 s), etc.
@@
Figure 2: Illustration of hidden point removal, without which the results would be messed up.
@@
Figure 3: Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh. Middle Left: An naive unprojection implementation. For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$. Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with incorrect color since it’s near an occlusion border, producing an artifact. Bottom Right: our proposed Non-Border-First Unprojection. Observing the correspondence between the occlusion border in 2D image space and the visibility border in UV space (both marked by sky-blue dashed lines), we prioritize non-border areas in UV space to avoid such artifacts.
3.1 Geometry Extraction: Point to Surface
@@
The first step in PointDreamer is to reconstruct an untextured mesh from the input point cloud. Since color information does not need to be considered in this step, many existing point-to-surface methods [39, 40, 28] can be employed. In our implementation, we directly adopt the state-of-the-art POCO [^1] https://github.com/valeoai/POCO [4] for its high performance. We conduct experiments comparing different geometry reconstruction methods in Section 4 and our supplementary file.
3.2 Projection: 3D to 2D